A Bibliography of Supercomputing ’92

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

25 October 2010 Version 1.04

Abstract

This bibliography records articles presented at the Supercomputing ’92 conference.

Title word cross-reference

N [WS92].

-body [WS92].

16-20 [IEE92]. 16se [CS92].

3D [PRZ92].

5 [PTCF92].

800 [BKW92].

’92 [IEE92]. access [AP92]. Achieving [GMS^{+}92]. affinity [ML92]. agents [Hau92]. algebraic [KSH^{+}92]. algorithm [AGZ92, BO92, CS92, MRX92]. algorithms [CK92, CR92, SA92, TJ92, vPMF92]. Alliant [BKW92]. allocation [HZ92]. alloys [FHP92]. analysis [CR92, GK92, MY92]. application [Kry92]. Applications [CR92, JP92, NB92]. approach [CR92]. architectures [MDSV92, MXKN92, vPMF92]. array [GMS^{+}92]. arrays [cC92]. artificial [WWHH92]. assets [CHW92]. Astrophysical [WS92]. award [Hau92].

B [CS92]. B-spline [CS92]. based [FVS92, MCG^{+}92, NNIK92, Rup92, WSG92]. batch [AZ92]. behavior [TJ92]. Benchmark [NN92, BBDS92, TS92]. benefits [AVAM92]. binary [FHP92]. bit [LLJ92]. bit-controlled [LLJ92]. blockability [CK92]. blocked [vPMF92]. body [WS92]. buffered [MY92].

Cache [FVS92, Mic92, NB92, Tay92]. caches [SH92, TJ92]. calculations [CPS92, PRZ92, Zit92]. CAMPUS [BKW92]. CAMPUS/800 [BKW92]. cancer [Hau92]. case [cC92, LSW92]. chains [NRS92]. channel [Rup92]. characterization [NN92]. Characterizing [TJ92]. chemical [Zit92]. Cholesky [VNS92]. circuit [Rup92]. class [HZ92]. Classification [WWHH92]. clusters [BN92, Hau92]. CM [PTCF92]. CM-5 [PTCF92]. code [Lio92, LCD^{+}92, MCG^{+}92, YW92]. coherence [CV92, JE92]. coherent [Mic92, NB92]. communication [MXKN92, PTCF92]. comparison [HM92]. ComPaSS [MXKN92]. compilation [HHKT92]. compile [CR92]. compile-time [CR92]. Compiler [CK92, MCG^{+}92, WSG92]. complex [PRZ92]. computational [BN92, VLLY92]. computations [Tay92]. computer [GMS^{+}92, PBDO92]. computers [KK92, LCD^{+}92, Sus92]. computing [Hau92]. Concurrent [BGMZ92]. Connection [MYJ92]. consistency [FVS92]. contention [Num92]. continuation [YS92]. continuous [MRX92]. controlled [LLJ92]. convective [MYJ92]. copy [BKW92]. CRAY [CS92, BO92, YW92]. curve [CS92].

D [HHKT92]. data [AP92, TR92, WS92]. debugging [NM92]. delays [AP92]. Delta [CPS92]. design [Hau92, SMRL92]. designs [Tay92]. desired [YS92]. dimensional [VLLY92]. Dimensions [Sch92]. direct [KSH^{+}92]. directional [FHP92]. discharge [YW92]. discrete [MRX92]. discrete-continuous [MRX92]. discussion [BBF^{+}92]. dissection [PSWB92]. distributed [BGY92, CR92, HHKT92, JE92, Kry92, MDSV92, Sus92, TR92, VNS92, vPMF92]. distributed-memory [HHKT92]. distribution [AG92, HHC92]. driven [BO92, Sus92]. DSP [GMS^{+}92]. Dynamic [TR92, AP92, Mic92].

EA [CS92]. EA/ [CS92]. Effect [AP92]. effective [CV92]. Efficient [DS92, WL92, MXKN92]. elastic [PRZ92]. Electromagnetic [CPS92]. element [GK92]. energy [MRX92]. environment [JA92, PHX^{+}92]. equations [JP92]. Euler [MDSV92]. event [BO92]. event-driven [BO92]. exits [HHC92]. exploitation [JA92].

factorization [VNS92, vPMF92]. fast [PSWB92]. Fault [AG92]. Fault-tolerant [AG92]. fibre [Rup92]. field [Bai92]. file [BGMZ92]. files [LS92]. finite [GK92, MY92]. fitting [CS92]. flow [YS92]. flows [MYJ92]. fluid [LCD^{+}92]. Forefronts [Hau92]. formation [FHP92]. FORTRAN [BGMZ92, HHKT92]. freckle [FHP92]. fully [YW92]. fully-vectorized [YW92].

generation [NN92]. genome [WWHH92]. geology [PRZ92]. gigabit [Rup92]. gigabit/sec [Rup92]. glow [YW92]. graphs [AVAM92, GK92].

heterogeneity [AVAM92]. Heterogeneous [RSL92, NRS92]. hierarchical [FVS92, WS92, WL92]. hierarchical-ring-based [FVS92]. High [Kry92, AGZ92, BGMZ92, MCG^{+}92]. high-performance [MCG^{+}92]. highly [Lio92]. HiPPI [JA92]. hypercube [AG92, HZ92]. hypercube-like [HZ92].

Implementation

[MDSV92, BO92, PSWB92]. implications [Tay92]. improved [TS92]. incorrect [BBF^{+}92]. information [CHW92, Fei92]. instruction [SH92]. integrating [WSG92]. Intel [CPS92, LCD^{+}92]. interactive [AZ92]. interconnected [cC92]. interconnection [KSH^{+}92, MY92, WL92]. Interprocedural [HHKT92]. investigate [AVAM92]. IP [JA92]. irregular [PTCF92]. issues [Kry92].

Jade [RSL92].

Lambda [LLJ92]. LAN [Rup92]. languages [BBF^{+}92]. Large [PRZ92, Hau92, JP92]. left [VLLY92]. like [HZ92]. Linda [DS92]. linear [JP92]. list [CR92]. load [AG92]. logic [BO92]. Loop [HHC92, ML92]. loops [Ram92]. Low [BKW92]. LU [vPMF92].

M3S [SMRL92]. Machine [MYJ92]. machines [HHKT92]. management [AG92, PHX^{+}92, TR92]. Mapping [NB92, Sus92]. massively [AG92, JP92, LSW92]. matrix [AGZ92, Tay92, Zit92]. matrix-vector [AGZ92]. meaning [Sch92]. Memory [Num92, AP92, AZ92, BDGS92, cC92, CR92, HHKT92, JE92, ML92, MDSV92, SMRL92, Sus92, vPMF92]. message [BKW92, NM92]. message-passing [NM92]. methods [WSG92, YS92, Zit92]. metrics [HM92]. MIMD [HHKT92]. minimization [MRX92]. Minneapolis [IEE92]. Minnesota [IEE92]. Misleading [Bai92]. mixed [AZ92]. mixing [MYJ92]. Model [Sus92, VLLY92]. Model-driven [Sus92]. modeling [KSH^{+}92, YW92]. molecular [MRX92]. MP [CS92, BO92, YW92]. multigrid [Lio92]. Multiparameter [YS92]. multiple [HHC92]. multiplication [AGZ92]. multiported [SMRL92]. multiprocessor [AP92, BDGS92, NB92, SMRL92, WL92]. multiprocessors [CR92, FVS92, JE92, ML92, NN92, Num92]. multistage [MY92]. Multitasking [SA92]. Multithreaded [PBDO92].

NAS [BBDS92]. national [CHW92]. nested [PSWB92, Ram92]. network [Kry92]. networks [KSH^{+}92, MY92, NRS92]. neural [WWHH92]. Non [Ram92]. Non-unimodular [Ram92]. nonequilibrium [YW92]. novel [Hau92]. November [IEE92]. Numerical [FHP92, CK92, Hau92].

object [TR92]. onto [NB92, Sus92].

operations [BGMZ92]. Optimal [NM92]. optimization [CR92, SA92].

panel [BBF^{+}92]. Parallel [BO92, HM92, AG92, AVAM92, AZ92, BBDS92, DS92, Hau92, JP92, KK92, LSW92, LCD^{+}92, MDSV92, NM92, RSL92, Sus92, YW92]. Parallel-and-vector [BO92]. parallelization [vPMF92]. Parallelizing [Lio92, DCS92, WSG92]. particle [MYJ92]. partitionability [AP92]. partitioned [LS92]. partitioning [GK92]. passing [BKW92, NM92]. patterns [PTCF92]. Performance [LCD^{+}92, MY92, VNS92, AP92, AGZ92, Bai92, BGMZ92, CR92, GMS^{+}92, HM92, LSW92, MCG^{+}92, NN92]. permuter [LLJ92]. pharmaceutical [Hau92]. pipeline [NNIK92]. plasma [LCD^{+}92]. pointing [Mic92]. policy [CV92]. Politically [BBF^{+}92]. polymer [NRS92]. potential [AVAM92]. pre [AGZ92, VNS92]. pre-processing [AGZ92]. pre-scheduling [VNS92]. Prefetching [SH92]. Proceedings [IEE92]. process [Zit92]. processing [AGZ92, YW92]. Processor [HZ92, GMS^{+}92, ML92, NNIK92]. processors [LSW92]. program [HM92]. programming [DS92, RSL92]. programs [DCS92, NM92]. propagation [PRZ92]. Pseudo [NNIK92].

random [AVAM92]. re [AG92]. re-distribution [AG92]. realities [LSW92]. register [LS92, NNIK92]. register-windowed [NNIK92]. regular [PTCF92]. relaxation [Lio92]. renewable [PHX^{+}92]. replay [NM92]. research [WWHH92]. resources [PHX^{+}92]. results [BBDS92, GK92, KK92]. RF [YW92]. ring [FVS92]. role [BN92].

Scalability [GK92]. scalable [BDGS92, JE92, MXKN92, Mic92]. scale [cC92, Hau92, PRZ92]. scattering [CPS92]. Scheduling [AZ92, PTCF92, CR92, ML92, VNS92, WSG92]. scheme [JE92, Mic92]. schemes [CV92]. search [KK92]. sec [Rup92]. separation [Zit92]. services [MXKN92]. shared [AP92, AZ92, BDGS92, JE92, ML92, MDSV92, Num92, SMRL92]. shared-memory [JE92, ML92, SMRL92]. SIMD [KK92]. Simulation [MYJ92, BO92, FHP92, Hau92, NRS92]. simulations [WS92]. slot [LLJ92]. snoopy [JE92]. software [CV92]. solidification [FHP92]. Solution [JP92]. solver [MDSV92]. sorting [TS92]. space [SA92]. Sparse [Tay92, Zit92, AGZ92, JP92, TJ92, VNS92]. spectral [PSWB92]. speed [Kry92]. spline [CS92]. states [YS92]. Storage [CHW92]. strategies [GK92]. structures [SA92, TR92, WS92]. study [LS92, LSW92]. summary [GK92, KK92]. super [GMS^{+}92]. supercomputer [AZ92, BO92, SH92, TS92]. supercomputers [HZ92, Zit92]. Supercomputing [IEE92, MWC^{+}92, Bai92, BBF^{+}92]. Superconcurrent [NRS92]. superscalar [MCG^{+}92, NNIK92]. superscalar-based [MCG^{+}92]. switch [Rup92]. system [AP92, DCS92, Mic92]. systems [AG92, AVAM92, CHW92, JP92, MCG^{+}92, PBDO92, WL92, WWHH92].

task [AG92, AVAM92]. TCP [JA92]. TCP/IP [JA92]. Techniques [WSG92]. theory [KSH^{+}92]. thin [VLLY92]. thin-walled [VLLY92]. three [VLLY92]. three-dimensional [VLLY92]. time [CR92, LLJ92]. tolerant [AG92]. Touchstone [CPS92]. tracing [NM92]. tracking [YS92]. transformations [MCG^{+}92, Ram92, WSG92]. treatment [Hau92]. tree [KK92, WS92]. turbulent [MYJ92].

uncertain [PHX^{+}92]. unimodular

[Ram92]. Unstructured [KK92, MDSV92]. Using [AVAM92, ML92, AGZ92, WS92].

validation [HM92]. vector

[AGZ92, BO92, LS92, NNIK92, Num92]. Vectorized [CS92, Lio92, YW92]. ventricle [VLLY92]. Virtual [Fei92, BGY92]. Visualization [PHX^{+}92, DCS92, Kry92]. visualizing [Fei92].

wafer [cC92]. wafer-scale [cC92]. walled [VLLY92]. wave [PRZ92]. Willow [BDGS92]. windowed [NNIK92]. windtunnel [BGY92]. workload [AZ92, NN92]. world [MWC^{+}92]. worlds [Fei92]. write [CV92].

X [CS92, YW92]. X-MP [CS92, YW92].

Y-MP [BO92].

zebra [Lio92].

References

Ahmad:1992:FTM

[AG92] I. Ahmad and A. Ghafoor. Faulttolerant task management and load re-distribution on massively parallel hypercube systems. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 750– 759. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Agarwal:1992:HPA

[AGZ92] R. C. Agarwal, F. G. Gustavson, and M. Zubair. A high performance algorithm using pre-processing for the sparse matrix-vector multiplication. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 32– 41. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Abraham:1992:EDA

[AP92] S. Abraham and K. Padmanabhan. Effect of data access delays and system partitionability on the dynamic performance of a shared memory multiprocessor. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 674– 682. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Almeida:1992:URT

[AVAM92] V. A. F. Almeida, I. M. M. Vasconcelos, J. N. C. Arabe, and D. A. Menasce. Using random task graphs to investigate the potential benefits of heterogeneity in parallel systems. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 683– 691. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Ashok:1992:SMI

[AZ92] I. Ashok and J. Zahorjan. Scheduling a mixed interactive and batch workload on a parallel, shared memory supercomputer. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 616– 625. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Bailey:1992:MPS

[Bai92] D. H. Bailey. Misleading performance in the supercomputing field. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 155–158. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Bailey:1992:NPB

[BBDS92] D. H. Bailey, E. Barszcz, L. Dagum, and H. D. Simon. NAS parallel benchmark results. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 386– 393. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Bohm:1992:PIL

[BBF^{+}92] W. Bohm, J. C. Browne, D. W. Forslund, A. Goforth, J. R. Mc-Graw, and K. S. Smith. Politically incorrect languages for supercomputing — a panel discussion. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 704–706. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Bennett:1992:WSS

[BDGS92] J. K. Bennett, S. Dwarkadas, J. Greenwood, and E. Speight. Willow: a scalable shared memory multiprocessor. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 336– 345. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Brezany:1992:CFO

[BGMZ92] P. Brezany, M. Gerndt, P. Mehrotra, and H. Zima. Concurrent file operations in a high performance FORTRAN. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 230–237. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Bryson:1992:DVW

[BGY92] S. Bryson and M. Gerald-Yamasaki. The distributed virtual windtunnel. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 275–284. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Burns:1992:LCM

[BKW92] C. M. Burns, R. H. Kuhn, and E. J. Werme. Low copy message passing on the Alliant CAM-PUS/800. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 760–769. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Brooks:1992:RCC

[BN92] E. D. Brooks, III and T. Nash. The role of computational clusters. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 162–164. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Bataineh:1992:PIE

[BO92] A. Bataineh and F. Ozguner. Parallel-and-vector implementation of the event-driven logic simulation algorithm on the Cray Y-MP supercomputer. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 444– 452. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-

Chiueh:1992:CWI

[cC92] T. c. Chiueh. A case for waferscale interconnected memory arrays. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 468–477. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Coyne:1992:SSN

[CHW92] R. A. Coyne, H. Hulen, and R. Watson. Storage systems for national information assets. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 626– 633. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Carr:1992:CBN

[CK92] S. Carr and K. Kennedy. Compiler blockability of numerical algorithms. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 114–124. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Cwik:1992:ESC

[CPS92] T. Cwik, J. Patterson, and D. Scott. Electromagnetic scattering calculations on the Intel Touchstone Delta. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 538–542. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Chung:1992:APA

[CR92] Yeh-Ching Chung and S. Ranka. Applications and performance analysis of a compile-time optimization approach for list scheduling algorithms on distributed memory multiprocessors. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 512–521. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Chung:1992:VAB

[CS92] Kuo-Liang Chung and Liang-Jzer Shen. Vectorized algorithm for B-spline curve fitting on CRAY X-MP EA/ 16se. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 166–169. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Chen:1992:EWP

[CV92] Y.-C. Chen and A. V. Veidenbaum. An effective write policy for software coherence schemes. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 661– 672. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Dow:1992:VSP

[DCS92] C.-R. Dow, S.-K. Chang, and M. I. Soffa. A visualization system for parallelizing programs. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 194– 203. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Deshpande:1992:EPP

[DS92] A. Deshpande and M. Schultz. Efficient parallel programming with Linda. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 238–244. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Feiner:1992:VWV

[Fei92] S. Feiner. Virtual worlds for visualizing information. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 638– 639. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title:

Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Felicelli:1992:NSF

[FHP92] S. D. Felicelli, J. C. Heinrich, and D. R. Poirier. Numerical simulation of freckle formation in directional solidification of binary alloys. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 433–441. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Farkas:1992:CCH

[FVS92] K. Farkas, Z. Vranesic, and M. Stumm. Cache consistency in hierarchical-ring-based multiprocessors. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 348–357. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Grama:1992:SAP

[GK92] A. Y. Grama and V. Kumar. Scalability analysis of partitioning strategies for finite element graphs: a summary of results. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 83– 92. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Gunzinger:1992:ASC

[GMS^{+}92] A. Gunzinger, U. A. Muller, W. Scott, B. Baumle, P. Kohler, H. R. vonder Muhll, F. Muller-Plathe, W. F. van Gunsteren, and W. Guggenbuhl. Achieving super computer performance with a DSP array processor. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 543– 550. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Hausheer:1992:FLS

[Hau92] F. H. Hausheer. Forefronts in large scale computing award: numerical simulation, parallel clusters, and the design of novel pharmaceutical agents for cancer treatment. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 636–637. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Hsieh:1992:LDM

[HHC92] B.-M. Hsieh, M. Hind, and R. Cytron. Loop distribution with multiple exits. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 204– 213. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Hall:1992:ICF

[HHKT92] M. W. Hall, S. Hiranandani, K. Kennedy, and C.-W. Tseng. Interprocedural compilation of Fortran D for MIMD distributedmemory machines. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 522– 534. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Hollingsworth:1992:PPP

[HM92] J. K. Hollingsworth and B. P. Miller. Parallel program performance metrics: a comparison and validation. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 4–13. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Haravu:1992:PAC

[HZ92] N. G. Haravu and S. G. Ziavras. Processor allocation for a class of hypercube-like supercomputers. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 740–749. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

IEEE:1992:PSM

[IEE92] IEEE Computer Society. Technical Committee on Computer Architecture, editor. Proceedings, Supercomputing ’92: Minneapolis, Minnesota, November 16-20, 1992. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Jung:1992:HET

[JA92] C. C.-D. Jung and D. Advani.

HiPPI exploitation in TCP/IP environment. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 610–615. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Jou:1992:SSC

[JE92] T.-S. Jou and R. Enbody. A scalable snoopy coherence scheme on distributed shared-memory multiprocessors. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 652–660. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Jones:1992:SLS

[JP92] M. T. Jones and P. E. Plassmann. Solution of large, sparse systems of linear equations in massively parallel applications. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 551– 560. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Karypis:1992:UTS

[KK92] G. Karypis and V. Kumar. Unstructured tree search on SIMD parallel computers: a summary of results. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 453–462. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Krystynak:1992:HSN

[Kry92] J. Krystynak. High speed network issues in a distributed visualization application. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 267– 274. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Kaushik:1992:ATM

[KSH^{+}92] S. D. Kaushik, S. Sharma, C.- H. Huang, J. R. Johnson, R. W. Johnson, and P. Sadayappan. An algebraic theory for modeling direct interconnection networks. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 488– 497. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title:

Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Lynch:1992:PPF

[LCD^{+}92] V. E. Lynch, B. A. Carreras, J. B. Drake, J. N. Leboeuf, and P. Liewer. Performance of a plasma fluid code on the Intel parallel computers. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 286– 293. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Lioen:1992:PHV

[Lio92]  W. M. Lioen.  Parallelizing

a highly vectorized multigrid code with zebra relaxation. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 180– 189. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Lee:1992:LTS

[LLJ92] D. Lee, K. Y. Lee, and K. F.

Jordan. The lambda time slot permuter: a bit-controlled time slot permuter. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 728–737. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Lee:1992:SPV

[LS92] C. G. Lee and J. E. Smith. A study of partitioned vector register files. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 94–103. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Lubeck:1992:PRM

[LSW92] O. M. Lubeck, M. L. Simmons, and H. J. Wasserman. The performance realities of massively parallel processors: a case study. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 403– 412. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Mahlke:1992:CCT

[MCG^{+}92] S. A. Mahlke, W. Y. Chen, J. C. Gyllenhaal, W. W. Hwu, P. P. Chang, and T. Kiyohara. Compiler code transformations for superscalar-based high-performance systems. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 808– 817. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Mavriplis:1992:IPU

[MDSV92] D. J. Mavriplis, R. Das, J. Saltz, and R. E. Vermeland. Implementation of a parallel unstructured Euler solver on shared and distributed memory architectures. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 132– 141. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Michael:1992:SCC

[Mic92] W. Michael. A scalable coherent cache system with a dynamic pointing scheme. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 358–367. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Markatos:1992:UPA

[ML92] E. P. Markatos and T. J.

LeBlanc. Using processor affinity in loop scheduling on sharedmemory multiprocessors. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 104– 113. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Maier:1992:DAM

[MRX92] R. S. Maier, J. B. Rosen, and G. L. Xue. A discrete-continuous algorithm for molecular energy minimization. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 778–786. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Malony:1992:SAW

[MWC^{+}92] A. D. Malony, D. J. Wallace, S.- C. Chang, M. A. McRobbie, V. P. Bhatkar, and K. Oliveira. Supercomputing around the world. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 126– 129. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

McKinley:1992:CEC

[MXKN92] P. K. McKinley, H. Xu, E. T. Kalns, and L. M. Ni. ComPaSS: efficient communication services for scalable architectures. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 478– 487. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Mun:1992:PAF

[MY92] Y. Mun and H. Y. Youn. Performance analysis of finite buffered multistage interconnection networks. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 718–727. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Malevsky:1992:SPM

[MYJ92] A. V. Malevsky, D. A. Yuen, and K. E. Jordan. Simulation of particle mixing by turbulent convective flows on the connection machine. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 294–300. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Nanda:1992:MAC

[NB92] A. K. Nanda and L. N. Bhuyan. Mapping applications onto a cache coherent multiprocessor. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 368– 377. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Netzer:1992:OTR

[NM92] R. H. B. Netzer and B. P. Miller. Optimal tracing and replay for debugging message-passing parallel programs. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 502–511. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Nanda:1992:BWG

[NN92] A. Nanda and L. M. Ni. Benchmark workload generation and performance characterization of multiprocessors. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 20–29. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Nakazawa:1992:PVP

[NNIK92] K. Nakazawa, H. Nakamura, H. Imori, and S. Kawabe. Pseudo vector processor based on register-windowed superscalar pipeline. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 642–651. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Nakanishi:1992:SSP

[NRS92] H. Nakanishi, V. Rego, and V. Sunderam. Superconcurrent simulation of polymer chains on heterogeneous networks. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 561– 569. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Numrich:1992:MCS

[Num92] R. W. Numrich. Memory contention for shared memory vector multiprocessors. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 316–325. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Papadopoulos:1992:MCS

[PBDO92] G. M. Papadopoulos, A. P. W. Bohm, A. T. Dahbura, and R. R. Oldehoeft. Multithreaded computer systems. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 772–775. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Pratico:1992:VMR

[PHX^{+}92] C. J. Pratico, F. B. Hanson, H. H.

Xu, D. J. Jarvis, and M. S. Vetter. Visualization for the management of renewable resources in an uncertain environment. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 258– 266. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Pereyra:1992:LSC

[PRZ92] V. Pereyra, E. Richardson, and S. E. Zarantonello. Large scale calculations of 3D elastic wave propagation in a complex geology. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 301–309. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Pothen:1992:TFI

[PSWB92] A. Pothen, H. D. Simon, L. Wang, and S. T. Barnard. Towards a fast implementation of spectral nested dissection. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 42– 51. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Ponnusamy:1992:SRI

[PTCF92] R. Ponnusamy, R. Thakur, A. Choudhary, and G. Fox. Scheduling regular and irregular communication patterns on the CM-5. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 394–402. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Ramanujam:1992:NTN

[Ram92] J. Ramanujam. Non-unimodular transformations of nested loops. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 214– 223. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Rinard:1992:HPP

[RSL92] M. C. Rinard, D. J. Scales, and M. S. Lam. Heterogeneous parallel programming in Jade. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 245– 256. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Rupert:1992:GSF

[Rup92] P. R. Rupert. A gigabit/sec fibre channel circuit switch based LAN. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], page 154. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title:

Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Saleh:1992:MAO

[SA92] A. Saleh and H. Adeli. Multitasking algorithms for optimization of space structures. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 64– 72. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Schutze:1992:DM

[Sch92] H. Schutze. Dimensions of meaning. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 787–796. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Smith:1992:PSI

[SH92] J. E. Smith and W.-C. Hsu. Prefetching in supercomputer instruction caches. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 588–597. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Sainrat:1992:DMM

[SMRL92] P. Sainrat, A. Mzoughi, C. Rochange, and D. Litaize. The design of the M3S: a multiported sharedmemory multiprocessor. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 326– 335. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Sussman:1992:MMD

[Sus92] A. Sussman. Model-driven mapping onto distributed memory parallel computers. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 818– 829. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Taylor:1992:SMC

[Tay92] V. E. Taylor. Sparse matrix computations: implications for cache designs. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 598–607. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Temam:1992:CBS

[TJ92] O. Temam and W. Jalby. Characterizing the behavior of sparse algorithms on caches. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 578– 587. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Totty:1992:DOM

[TR92] B. K. Totty and D. A. Reed.

Dynamic object management for distributed data structures. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 692– 701. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Thearling:1992:ISS

[TS92] K. Thearling and S. Smith. An improved supercomputer sorting benchmark. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 14–19. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Vesier:1992:TCM

[VLLY92] C. C. Vesier, J. D. Lemmon, Jr., R. A. Levine, and A. P. Yoganathan. A three-dimensional computational model of a thinwalled left ventricle. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 73– 82. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Venugopal:1992:PDS

[VNS92] S. Venugopal, V. K. Naik, and J. Saltz. Performance of distributed sparse Cholesky factorization with pre-scheduling. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 52– 61. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

vonLaszewski:1992:PBL

[vPMF92] G. von Laszewski, M. Parashar, A. G. Mohamed, and G. C. Fox. On the parallelization of blocked LU factorization algorithms on distributed memory architectures. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 170–179. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Wei:1992:EHI

[WL92] S. Wei and S. Levy. Efficient hierarchical interconnection for multiprocessor systems. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 708– 717. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Warren:1992:ANS

[WS92] Michael S. Warren and John K.

Salmon. Astrophysical Nbody simulations using hierarchical tree data structures. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 570– 576. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. URL ftp://

ftp.cacr.caltech.edu/nbody/

sc92.ps.Z. 1992 Gordon Bell Prize winner.

Watts:1992:TIP

[WSG92] T. M. Watts, M. L. Soffa, and R. Gupta. Techniques for integrating parallelizing transformations and compiler based scheduling methods. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 830–839. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Wu:1992:CAN

[WWHH92] C. H. Wu, G. M. Whitson, C.-T. Hsiao, and C.-F. Huang. Classification artificial neural systems for genome research. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 797– 803. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Yeckel:1992:MCM

[YS92] A. Yeckel and L. E. Scriven. Multiparameter continuation methods for tracking desired flow states. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 142–151. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Young:1992:FCN

[YW92] F. F. Young and C.-H. Wu. A fully-vectorized code for nonequilibrium RF glow discharge modeling and its parallel processing on a Cray X-MP. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 424–432. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Zitney:1992:SMM

[Zit92] S. E. Zitney. Sparse matrix methods for chemical process separation calculations on supercomputers. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 414–423. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.